OpenStudy (anonymous):

quadratic equation of 3x-2xsquared=7

OpenStudy (eric_d):

3x-2x^2=7 -2x^2-7=0 factorise it ( )( )=0

OpenStudy (anonymous):

oh okay .. thanks :) <3

OpenStudy (eric_d):

np